### v1

# On discrete Sobolev inequalities for nonconforming finite elements under a semi-regular mesh condition

Abstract
Mathematics Subject Classification (2020) 65N30 (primary); 65N15, 65N12, 46E35.
1 Introduction
This paper introduces discrete Sobolev inequalities applicable to nonconforming finite elements on semi-regular (anisotropic) meshes. Specifically, for the Crouzeix–Raviart (CR) finite element space and its discontinuous counterpart, we establish an $L^{q}-L^{p}$ ($q\leq p$) bound. Notably, the constants associated with this bound are influenced solely by the domain and the semi-regular parameter, remaining unaffected by the angles and aspect ratios of simplices. This extends the classical shape-regular theory (e.g. [2, 16]) to anisotropic partitions and provides a robust tool for stability and error analysis.
Let $\Omega\subset\mathbb{R}^{d}$, $d\in\{2,3\}$ be a bounded polyhedral domain. The Poisson problem is to find $u:\Omega\to\mathbb{R}$ such that
$\displaystyle-\varDelta u=f\quad\text{in $\Omega$},\quad u=0\quad\text{on $\partial\Omega$},$ (1.1)
where $f\in L^{2}(\Omega)$ is a given function. In [14, 7, 11], we considered the CR finite element type discretisation as follows. Find $u_{h}\in V_{h}$ such that
$\displaystyle a_{h}^{CR}(u_{h},v_{h}):=\int_{\Omega}\nabla_{h}u_{h}\cdot\nabla_{h}v_{h}dx+\sum_{F\in\mathcal{F}_{h}}\kappa_{F}\int_{F}\Pi_{F}^{0}[\![u_{h}]\!]\Pi_{F}^{0}[\![v_{h}]\!]ds$ $=\int_{\Omega}fv_{h}dx\quad\forall v_{h}\in V_{h},$ (1.2)
where $V_{h}$ is the CR finite element space defined in Section 2.3, $\kappa_{F}$ is a penalty parameter defined in [7, Section 2.4] and $\Pi_{F}^{0}$ is the $L^{2}$-projection defined in Section 2.4. We define an energy norm as
$\displaystyle|v_{h}|_{E}:=a_{h}^{CR}(v_{h},v_{h})^{\frac{1}{2}},\quad v_{h}\in V_{h}.$
Stability in the finite element method is the property that the norm of the discrete solution is controlled by a uniform constant, independent of the mesh and parameters, in the norm of the right-hand side (external force) data, which guarantees the reliability of the numerical solution. Substituting the finite element solution $u_{h}\in V_{h}$ for $v_{h}$ on the right-hand side of (1.2) yields
$\displaystyle|u_{h}|_{E}^{2}=\int_{\Omega}fv_{h}dx\leq\|f\|_{L^{2}(\Omega)}\|u_{h}\|_{L^{2}(\Omega)}.$
If the discrete Poincaré inequality:
$\displaystyle\|u_{h}\|_{L^{2}(\Omega)}\leq C^{P}_{dc}|u_{h}|_{E}$
holds, the following stability estimate is obtained.
$\displaystyle|u_{h}|_{E}\leq C^{P}_{dc}\|f\|_{L^{2}(\Omega)}.$
For piecewise $H^{1}$ functions on general partitions, the Poincaré–Friedrichs inequalities are valid with constants that depend solely on the shape-regularity of the partition, without necessitating quasi-uniformity. In two dimensions ($2D$), shape-regularity can be assessed through triangulations that adhere to certain criteria, such as a minimum-angle parameter. In three dimensions ($3D$), it is determined by geometric parameters in conjunction with a uniform face-angle bound, as detailed in [2].
Demonstrating the discrete Sobolev (Poincaré) inequality on anisotropic meshes presents significant challenges. Recent research has made notable progress in this area. The study in [7] established the discrete Poincaré inequality for the anisotropic weakly over-penalised symmetric interior penalty method. Following this, [9] extended these findings to a hybrid version of the method. Furthermore, [10] introduced the discrete Sobolev inequality on anisotropic meshes, specifically for the CR finite element method. The discrete Poincaré inequality as applied to Nitsche’s method on anisotropic meshes is detailed in [11]. These investigations utilise proofs based on the duality problem, necessitating the assumption of a convex domain $\Omega$ to substantiate the inequalities. This paper aims to give a new anisotropic discrete Sobolev inequality that removes the assumption that the domain is convex. Our results extend the classical shape-regular framework of Brenner (2003) to semi-regular meshes. For $q\mathop{\textgreater}p$, an $L^{p}-L^{2}$ estimate is available under a weak elliptic regularity (see Section 6 and [10, Lemma 4]).
Section 2 establishes the notation and anisotropic preliminaries. Section 3 introduces semi-regular meshes and the Raviart–Thomas (RT) interpolation. Section 4 compiles the Bogovskiĭ operator, discrete integration-by-parts, and face estimates. Section 5 demonstrates the primary discrete $L^{q}-L^{p}$ inequality, while Section 6 provides concluding remarks.
Throughout this paper, generic constants $c$ depend only on $d$, $\Omega$, $p$ and the semi-regular parameter; they are independent of $h$ (defined later), angles and aspect ratios of simplices. These values vary across different contexts. The notation $\mathbb{R}_{+}$ denotes a set of positive real numbers.
2 Preliminaries
Cross-references. See Section 2.1 for the jumps/averages and boundary treatment, Section 2.3 for the definitions of the CR/discontinuous-CR spaces and Section 2.4 for the discrete norms.
2.1 Mesh, Mesh faces, jumps and average
Let $\mathbb{T}_{h}=\{T\}$ be a simplicial mesh of $\overline{\Omega}$ composed of closed $d$-simplices:
$\displaystyle\overline{\Omega}=\bigcup_{T\in\mathbb{T}_{h}}T,$
where $h:=\max_{T\in\mathbb{T}_{h}}h_{T}$ and $h_{T}:=\mathop{\mathrm{diam}}(T)$. For simplicity, we assume that $\mathbb{T}_{h}$ is conformal; that is, $\mathbb{T}_{h}$ is a simplicial mesh of $\overline{\Omega}$ without hanging nodes. Throughout we assume $h\leq 1$.
Let $\mathcal{F}_{h}^{i}$ be the set of interior faces, and $\mathcal{F}_{h}^{\partial}$ be the set of faces on the boundary $\partial\Omega$. We set $\mathcal{F}_{h}:=\mathcal{F}_{h}^{i}\cup\mathcal{F}_{h}^{\partial}$. For any $F\in\mathcal{F}_{h}$, we define the unit normal $\bm{n}_{F}$ to $F$ as follows: (i) If $F\in\mathcal{F}_{h}^{i}$ with $F=T_{\natural}\cap T_{\sharp}$, $T_{\natural},T_{\sharp}\in\mathbb{T}_{h}$, $\natural>\sharp$, let $\bm{n}_{F}$ be the unit normal vector from $T_{\natural}$ to $T_{\sharp}$. (ii) If $F\in\mathcal{F}_{h}^{\partial}$, $\bm{n}_{F}$ is the unit outward normal $\bm{n}$ to $\partial\Omega$.
Let $p\in[1,\infty)$ and $s\mathop{\textgreater}0$ be a positive real number. We define a broken (piecewise) Sobolev space as
$\displaystyle W^{s,p}(\mathbb{T}_{h}):=\{v\in L^{p}(\Omega):\ v|_{T}\in W^{s,p}(T)\ \forall T\in\mathbb{T}_{h}\}$
with a norm
$\displaystyle|v|_{W^{s,p}(\mathbb{T}_{h})}$ $:=\left(\sum_{T\in\mathbb{T}_{h}}|v|^{p}_{W^{s,p}(T)}\right)^{\frac{1}{p}}\quad\text{if $p\in[1,\infty)$}.$
Especially, we write $H^{s}(\mathbb{T}_{h}):=W^{s,2}(\mathbb{T}_{h})$ with a norm
$\displaystyle|v|_{H^{s}(\mathbb{T}_{h})}$ $:=\left(\sum_{T\in\mathbb{T}_{h}}|v|^{2}_{H^{s}(T)}\right)^{\frac{1}{2}}\quad v\in H^{s}(\mathbb{T}_{h}).$
Let $p\in[1,\infty)$. Let $\varphi\in W^{1,p}(\mathbb{T}_{h})$. Suppose that $F\in\mathcal{F}_{h}^{i}$ with $F=T_{+}\cap T_{-}$, $T_{+},T_{-}\in\mathbb{T}_{h}$, +>+>-. We set $\varphi_{+}:=\varphi{|_{T_{+}}}$ and $\varphi_{-}:=\varphi{|_{T_{-}}}$. We set two nonnegative real numbers $\omega_{T_{+},F}$ and $\omega_{T_{-},F}$ such that
$\displaystyle\omega_{T_{+},F}+\omega_{T_{-},F}=1.$
The jump and skew-weighted averages of $\varphi$ across $F$ are then defined as
$\displaystyle[\![\varphi]\!]:=[\![\varphi]\!]_{F}:=\varphi_{+}-\varphi_{-},\quad\{\!\{\varphi\}\!\}_{\overline{\omega}}:=\{\!\{\varphi\}\!\}_{\overline{\omega},F}:=\omega_{T_{-},F}\varphi_{+}+\omega_{T_{+},F}\varphi_{-},\quad+>-.$
For a boundary face $F\in\mathcal{F}_{h}^{\partial}$ with $F=\partial T\cap\partial\Omega$, $\varphi]\!]_{F}:=\varphi|_{T}$ and $\{\!\{\varphi\}\!\}_{\overline{\omega}}:=\varphi|_{T}$. For any ${\bm{v}}\in W^{1,p}(\mathbb{T}_{h})^{d}$, the notation
$[\![{\bm{v}}\cdot{\bm{n}}]\!]:=[\![{\bm{v}}\cdot{\bm{n}}]\!]_{F}:={\bm{v}_{+}}\cdot{\bm{n}_{F}}-{\bm{v}_{-}}\cdot{\bm{n}_{F}},\quad+>-,$
$\{\!\{{\bm{v}}\}\!\}_{\omega}:=\{\!\{{\bm{v}}\}\!\}_{\omega,F}:=\omega_{T_{+},F}{\bm{v}_{+}}+\omega_{T_{-},F}{\bm{v}_{-}},\quad+>$-
for the jump in the normal component and the weighted average of $\bm{v}$}. For any ${\bm{v}}\in W^{1,p}(\mathbb{T}_{h})^{d}$ and $\varphi\in W^{1,p}(\mathbb{T}_{h})$, we have that
$\displaystyle[\![({\bm{v}}\varphi)\cdot{\bm{n}}]\!]_{F}$ $=\{\!\{{\bm{v}}\}\!\}_{\omega,F}\cdot{\bm{n}_{F}}[\![\varphi]\!]_{F}+[\![{\bm{v}}\cdot{\bm{n}}]\!]_{F}\{\!\{\varphi\}\!\}_{\overline{\omega},F}.$
2.2 Trace inequality
The trace inequality on anisotropic meshes discussed herein is of considerable importance to this study. The proof of this inequality is documented in several references. In this context, we adhere to the approach outlined by Ern and Guermond [4, Lemma 12.15]. It is noteworthy that, although Lemma 12.15 in [4] stipulates a shape-regular mesh condition, the condition is easily violated. For a simplex $T\subset\mathbb{R}^{d}$, let $\mathcal{F}_{T}$ be the collection of the faces of $T$. Let $|\cdot|_{d}$ denote the $d$-dimensional Hausdorff measure.
Lemma 2.1 (Trace inequality).
Let $p\in[1,\infty]$. Let $T\subset\mathbb{R}^{d}$ be a simplex. There exists a positive constant $c$ such that for any $\bm{v}=(v^{(1)},\ldots,v^{(d)})^{\top}\in W^{1,p}(T)$, $F\in\mathcal{F}_{T}$, and $h$,
$\displaystyle\|\bm{v}\|_{L^{p}(F)^{d}}\leq c\ell_{T,F}^{-\frac{1}{p}}\left(\|\bm{v}\|_{L^{p}(T)^{d}}+h_{T}^{\frac{1}{p}}\|\bm{v}\|_{L^{p}(T)^{d}}^{1-\frac{1}{p}}|\bm{v}|_{W^{1,p}(T)^{d}}^{\frac{1}{p}}\right),$ (2.1)
where $\ell_{T,F}:=\frac{d!|T|_{d}}{|F|_{d-1}}$ denotes the distance of the vertex of $T$ opposite to $F$ to the face.
Proof.
Let $v\in W^{1,p}(T)$. Let $\bm{P}_{F}$ be the vertex of $T$ opposite to $F$. By the same argument of [4, Lemma 12.15], together with the fact that $|\bm{x}-\bm{P}_{F}|\leq h_{T}$ and $\frac{|F|_{d-1}}{|T|_{d}}=\frac{d!}{\ell_{T,F}}$, it holds that for $i\in\{1,\ldots,d\}$,
$\displaystyle\|v^{(i)}\|^{p}_{L^{p}(F)}$ $\leq\frac{d!}{\ell_{T,F}}\|v^{(i)}\|^{p}_{L^{p}(T)}+\frac{2(d-1)!}{\ell_{T,F}}h_{T}\|v^{(i)}\|^{p-1}_{L^{p}(T)}\|\bm{\nabla}v^{(i)}\|_{L^{p}(T)^{d}}.$
Using the Cauchy–Schwarz inequality, we obtain the target inequality together with Jensen’s inequality. ∎
Remark 2.2.
Because $|T|_{d}\approx h_{T}^{d}$ and $|F|_{d-1}\approx h_{T}^{d-1}$ on the shape-regular mesh, it holds that $\ell_{T,F}\approx h_{T}$. Then, the trace inequality (2.1) is given as
$\displaystyle\|v\|_{L^{p}(F)}\leq ch_{T}^{-\frac{1}{p}}\left(\|v\|_{L^{p}(T)}+h_{T}^{\frac{1}{p}}\|v\|_{L^{p}(T)}^{1-\frac{1}{p}}|v|_{W^{1,p}(T)}^{\frac{1}{p}}\right).$
2.3 Finite element spaces
For any $T\in\mathbb{T}_{h}$ and $F\in\mathcal{F}_{h}$, let $D\in\{T,F\}$. For $k\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$, $\mathbb{P}^{k}(D)$ is spanned by the restriction to $D$ of polynomials in $\mathbb{P}^{k}$, where $\mathbb{P}^{k}$ denotes the space of polynomials with a maximum of $k$ degrees.
For $k\in\mathbb{N}_{0}$, we define the standard discontinuous finite-element space as
$\displaystyle V_{h}^{DC(k)}$ $:=\left\{v_{h}\in{L^{\infty}(\Omega)};\ v_{h}|_{T}\in\mathbb{P}^{k}({T})\quad\forall T\in\mathbb{T}_{h}\right\}.$ (2.2)
Let $Ne$ be the number of elements included in the mesh $\mathbb{T}_{h}$. Thus, we write $\mathbb{T}_{h}=\{T_{j}\}_{j=1}^{Ne}$. We introduce a discontinuous CR finite element space.
Let the points $\{\bm{P}_{T_{j},1},\ldots,\bm{P}_{T_{j},d+1}\}$ be the vertices of the simplex $T_{j}\in\mathbb{T}_{h}$ for $j\in\{1,\ldots,Ne\}$. Let $F_{T_{j},i}$ be the face of $T_{j}$ opposite $\bm{P}_{T_{j},i}$ for $i\in\{1,\ldots,d+1\}$. We take a set ${\Sigma}_{T_{j}}:=\{{\chi}^{CR}_{T_{j},i}\}_{1\leq i\leq d+1}$ of linear forms with its components such that for any $q\in\mathbb{P}^{1}$.
$\displaystyle{\chi}^{CR}_{T_{j},i}({q}):=\frac{1}{|{F}_{T_{j},i}|_{d-1}}\int_{{F}_{T_{j},i}}{q}d{s}\quad\forall i\in\{1,\ldots,d+1\}.$ (2.3)
For each $j\in\{1,\ldots,Ne\}$, the triple $\{T_{j},\mathbb{P}^{1},\Sigma_{T_{j}}\}$ is a finite element. Using the barycentric coordinates $\{{\lambda}_{T_{j},i}\}_{i=1}^{d+1}:\mathbb{R}^{d}\to\mathbb{R}$ on the reference element, the nodal basis functions associated with the degrees of freedom by (2.3) are defined as
$\displaystyle{\theta}^{CR}_{T_{j},i}(\bm{x}):=d\left(\frac{1}{d}-{\lambda}_{T_{j},i}(\bm{x})\right)\quad\forall i\in\{1,\ldots,d+1\}.$ (2.4)
For $j\in\{1,\ldots,Ne\}$ and $i\in\{1,\ldots,d+1\}$, we define the function $\phi_{j(i)}$ as
$\displaystyle\phi_{j(i)}(\bm{x}):=\begin{cases}\theta_{T_{j},i}^{CR}(\bm{x}),\quad\text{$\bm{x}\in T_{j}$},\\ 0,\quad\text{$x\notin T_{j}$}.\end{cases}$ (2.5)
We define a discontinuous CR finite element space as
$\displaystyle V_{h}^{DCCR}$ $:=\left\{\sum_{j=1}^{Ne}\sum_{i=1}^{d+1}c_{j(i)}\phi_{j(i)};\ c_{j(i)}\in\mathbb{R},\ \forall i,j\right\}\subset V_{h}^{DC(1)}.$ (2.6)
Furthermore, we define standard CR finite element spaces as
$\displaystyle V_{h}^{CR}$ $:=\biggl{\{}\varphi_{h}\in V_{h}^{DCCR}:\ \int_{F}[\![\varphi_{h}]\!]ds=0\ \forall F\in\mathcal{F}_{h}^{i}\biggr{\}},$ (2.7)
$V_{h0}^{CR}$ $:=\biggl{\{}\varphi_{h}\in V_{h}^{CR}:\ \int_{F}\varphi_{h}ds=0\ \forall F\in\mathcal{F}_{h}^{\partial}\biggr{\}}.$ (2.8)
We define $V_{h}\in\{V_{h}^{DCCR},V_{h}^{CR},V_{h0}^{CR}\}$.
2.4 Norms
Let $F\in\mathcal{F}_{h}^{i}$ with $F=T_{+}\cap T_{-}$, $T_{+},T_{-}\in\mathbb{T}_{h}$, +>+>- be an interior face and $F\in\mathcal{F}_{h}^{\partial}$ with $F=\partial T_{\partial}\cap\partial\Omega$, $T_{\partial}\in\mathbb{T}_{h}$ a boundary face. Let $p\in[1,\infty)$. A choice for the weighted parameters is such that
$\displaystyle\omega_{T_{i},F}:=\frac{{\ell_{T_{i},F}^{\frac{p-1}{p}}}}{{\ell_{T_{+},F}^{\frac{p-1}{p}}}+{\ell_{T_{-},F}^{\frac{p-1}{p}}}},\quad i\in\{+,-\}.$ (2.9)
For the proof of the discrete Sobolev inequality, we will use the following parameter.
$\displaystyle\kappa_{p,F*}:=\begin{cases}\displaystyle\left({\ell_{T_{+},F}^{\frac{p-1}{p}}}+{\ell_{T_{-},F}^{\frac{p-1}{p}}}\right)^{-p}\quad\text{if $F\in\mathcal{F}_{h}^{i}$},\\ \displaystyle\ell_{T_{\partial},F}^{1-p}\quad\text{if $F\in\mathcal{F}_{h}^{\partial}$}.\end{cases}$ (2.10)
where $\ell_{T,F}$ is defined in Section 2.2. See [7, Section 2.4] for more information on this parameter.
Let $p\in[1,\infty)$ and $V_{h}\in\{V_{h}^{DCCR},V_{h}^{CR},V_{h0}^{CR}\}$. We define the following three norms for any $\varphi_{h}\in V_{h}$;
$\displaystyle|\varphi_{h}|_{p,V_{h}}$ $:=\left(|\varphi_{h}|_{W^{1,p}(\mathbb{T}_{h})}^{p}+|\varphi_{h}|_{p,J}^{p}\right)^{\frac{1}{p}}\text{ with }|\varphi_{h}|_{p,J}:=\left(\sum_{F\in\mathcal{F}_{h}}\kappa_{p,F*}\|\Pi_{F}^{0}[\![\varphi_{h}]\!]\|_{L^{p}(F)}^{p}\right)^{\frac{1}{p}},$
where for any $F\in\mathcal{F}_{h}$, we define the $L^{2}$-projection $\Pi_{F}^{0}:L^{2}(F)\to\mathbb{P}^{0}(F)$ as
$\displaystyle\int_{F}(\Pi_{F}^{0}\varphi-\varphi)ds=0\quad\forall\varphi\in L^{2}(F).$
We note that
$\displaystyle|\varphi_{h}|_{p,J}$ $=\left(\sum_{F\in\mathcal{F}_{h}^{\partial}}\kappa_{p,F*}\|\Pi_{F}^{0}[\![\varphi_{h}]\!]\|_{L^{p}(F)}^{p}\right)^{\frac{1}{p}}\quad\text{if $\varphi_{h}\in V_{h}^{CR}$},$
$|\varphi_{h}|_{p,J}$ $=0\quad\text{if $\varphi_{h}\in V_{h0}^{CR}$}.$
2.5 Reference elements
We first define the reference elements $\widehat{T}\subset\mathbb{R}^{d}$.
2.5.1 Two-dimensional case
Let $\widehat{T}\subset\mathbb{R}^{2}$ be a reference triangle with vertices $\hat{\bm{p}}_{1}:=(0,0){{}^{\top}}$, $\hat{\bm{p}}_{2}:=(1,0){{}^{\top}}$, and $\hat{\bm{p}}_{3}:=(0,1){{}^{\top}}$.
2.5.2 Three-dimensional case
In the three-dimensional case, we consider the following two cases: (i) and (ii); see Condition 2.4.
Let $\widehat{T}_{1}$ and $\widehat{T}_{2}$ be reference tetrahedra with the following vertices:
$\widehat{T}_{1}$ has vertices $\hat{\bm{p}}_{1}:=(0,0,0){{}^{\top}}$, $\hat{\bm{p}}_{2}:=(1,0,0){{}^{\top}}$, $\hat{\bm{p}}_{3}:=(0,1,0){{}^{\top}}$, and $\hat{\bm{p}}_{4}:=(0,0,1){{}^{\top}}$;
$\widehat{T}_{2}$ has vertices $\hat{\bm{p}}_{1}:=(0,0,0){{}^{\top}}$, $\hat{\bm{p}}_{2}:=(1,0,0){{}^{\top}}$, $\hat{\bm{p}}_{3}:=(1,1,0){{}^{\top}}$, and $\hat{\bm{p}}_{4}:=(0,0,1){{}^{\top}}$.
[…]
2.6 Two-step affine mapping
In anisotropic meshes, the mesh shape and element proportions are non-uniform, which directly affects the interpolation accuracy. Existing interpolation error estimates typically assume an even or regular mesh, which may overestimate the error if applied directly to anisotropic meshes. To remedy this, we proposed in [6, 8, 13, 15] a new strategy on anisotropic meshes.
To an affine simplex $T\subset\mathbb{R}^{d}$, we construct two affine mappings $\Phi_{\widetilde{T}}:\widehat{T}\to\widetilde{T}$ and $\Phi_{T}:\widetilde{T}\to T$. First, we define the affine mapping $\Phi_{\widetilde{T}}:\widehat{T}\to\widetilde{T}$ as
$\displaystyle\Phi_{\widetilde{T}}:\widehat{T}\ni\hat{\bm{x}}\mapsto\tilde{\bm{x}}:=\Phi_{\widetilde{T}}(\hat{\bm{x}}):={A}_{\widetilde{T}}\hat{\bm{x}}\in\widetilde{T}$, (2.11)
where ${A}_{\widetilde{T}}\in\mathbb{R}^{d\times d}$ is an invertible matrix. We then define the affine mapping $\Phi_{T}:\widetilde{T}\to T$ as follows:
$\displaystyle\Phi_{T}:\widetilde{T}\ni\tilde{\bm{x}}\mapsto\bm{x}:=\Phi_{T}(\tilde{\bm{x}}):={A}_{T}\tilde{\bm{x}}+\bm{b_{T}}\in T,$ (2.12)
where $\bm{b_{T}}\in\mathbb{R}^{d}$ is a vector and ${A}_{T}\in O(d)$ denotes the rotation and mirror-imaging matrix. We define the affine mapping $\Phi:\widehat{T}\to T$ as
$\displaystyle\Phi:={\Phi}_{T}\circ{\Phi}_{\widetilde{T}}:\widehat{T}\ni\hat{\bm{x}}\mapsto\bm{x}:=\Phi(\hat{\bm{x}})=({\Phi}_{T}\circ{\Phi}_{\widetilde{T}})(\hat{\bm{x}})={A}\hat{\bm{x}}+\bm{b_{T}}\in T,$
where ${A}:={A}_{T}{A}_{\widetilde{T}}\in\mathbb{R}^{d\times d}$.
2.6.1 Construct mapping $\Phi_{\widetilde{T}}:\widehat{T}\to\widetilde{T}$
We consider the affine mapping (2.11). We define the matrix ${A}_{\widetilde{T}}\in\mathbb{R}^{d\times d}$ as follows. We first define the diagonal matrix as
$\displaystyle\widehat{A}:=\mathop{\mathrm{diag}}(h_{1},\ldots,h_{d}),\quad h_{i}\in\mathbb{R}_{+}\quad\forall i.$ (2.13)
For $d=2$, we define the regular matrix $\widetilde{A}\in\mathbb{R}^{2\times 2}$ as
$\displaystyle\widetilde{A}:=\begin{pmatrix}1&s\\ 0&t\\ \end{pmatrix}$, (2.14)
with the parameters
$\displaystyle s^{2}+t^{2}=1,\quad t\mathop{\textgreater}0.$
For the reference element $\widehat{T}$, let $\mathfrak{T}^{(2)}$ be a family of triangles.
$\displaystyle\widetilde{T}$ $=\Phi_{\widetilde{T}}(\widehat{T})={A}_{\widetilde{T}}(\widehat{T}),\quad{A}_{\widetilde{T}}:=\widetilde{A}\widehat{A}$
with the vertices $\tilde{\bm{p}}_{1}:=(0,0)^{\top}$, $\tilde{\bm{p}}_{2}:=(h_{1},0)^{\top}$ and $\tilde{\bm{p}}_{3}:=(h_{2}s,h_{2}t)^{\top}$. Then, $h_{1}=|\tilde{\bm{p}}_{1}-\tilde{\bm{p}}_{2}|\mathop{\textgreater}0$ and $h_{2}=|\tilde{\bm{p}}_{1}-\tilde{\bm{p}}_{3}|\mathop{\textgreater}0$.
For $d=3$, we define the regular matrices $\widetilde{A}_{1},\widetilde{A}_{2}\in\mathbb{R}^{3\times 3}$ as follows:
$\displaystyle\widetilde{A}_{1}:=\begin{pmatrix}1&s_{1}&s_{21}\\ 0&t_{1}&s_{22}\\ 0&0&t_{2}\\ \end{pmatrix},\ \widetilde{A}_{2}:=\begin{pmatrix}1&-s_{1}&s_{21}\\ 0&t_{1}&s_{22}\\ 0&0&t_{2}\\ \end{pmatrix}$ (2.15)
with the parameters
$\displaystyle\begin{cases}s_{1}^{2}+t_{1}^{2}=1,\ s_{1}\mathop{\textgreater}0,\ t_{1}\mathop{\textgreater}0,\ h_{2}s_{1}\leq h_{1}/2,\\ s_{21}^{2}+s_{22}^{2}+t_{2}^{2}=1,\ t_{2}\mathop{\textgreater}0,\ h_{3}s_{21}\leq h_{1}/2.\end{cases}$
Therefore, we set $\widetilde{A}\in\{\widetilde{A}_{1},\widetilde{A}_{2}\}$. For the reference elements $\widehat{T}_{i}$, $i=1,2$, let $\mathfrak{T}_{i}^{(3)}$, $i=1,2$, be a family of tetrahedra.
$\displaystyle\widetilde{T}_{i}$ $=\Phi_{\widetilde{T}_{i}}(\widehat{T}_{i})={A}_{\widetilde{T}_{i}}(\widehat{T}_{i}),\quad{A}_{\widetilde{T}_{i}}:=\widetilde{A}_{i}\widehat{A},\quad i=1,2,$
with the vertices
$\tilde{\bm{p}}_{1}:=(0,0,0)^{\top},\ \tilde{\bm{p}}_{2}:=(h_{1},0,0)^{\top},\ \tilde{\bm{p}}_{4}:=(h_{3}s_{21},h_{3}s_{22},h_{3}t_{2})^{\top},$
$\begin{cases}\tilde{\bm{p}}_{3}:=(h_{2}s_{1},h_{2}t_{1},0)^{\top}\quad\text{for case (i)},\\ \tilde{\bm{p}}_{3}:=(h_{1}-h_{2}s_{1},h_{2}t_{1},0)^{\top}\quad\text{for case (ii)}.\end{cases}$
Subsequently, $h_{1}=|\tilde{\bm{p}}_{1}-\tilde{\bm{p}}_{2}|\mathop{\textgreater}0$, $h_{3}=|\tilde{\bm{p}}_{1}-\tilde{\bm{p}}_{4}|\mathop{\textgreater}0$, and
$\displaystyle h_{2}=\begin{cases}|\tilde{\bm{p}}_{1}-\tilde{\bm{p}}_{3}|\mathop{\textgreater}0\quad\text{for case (i)},\\ |\tilde{\bm{p}}_{2}-\tilde{\bm{p}}_{3}|\mathop{\textgreater}0\quad\text{for case (ii)}.\end{cases}$
2.6.2 Construct mapping $\Phi_{T}:\widetilde{T}\to T$
We determine the affine mapping (2.12) as follows. Let ${T}\in\mathbb{T}_{h}$ have vertices ${p}_{i}$ ($i=1,\ldots,d+1$). Let $b_{T}\in\mathbb{R}^{d}$ be the vector and ${A}_{T}\in O(d)$ be the rotation and mirror imaging matrix such that
$\displaystyle\bm{p}_{i}=\Phi_{T}(\tilde{\bm{p}}_{i})={A}_{T}\tilde{\bm{p}}_{i}+\bm{b_{T}},\quad i\in\{1,\ldots,d+1\},$
where vertices $\bm{p}_{i}$ ($i=1,\ldots,d+1$) satisfy the following conditions:
Condition 2.3 (Case in which $d=2$).
Let ${T}\in\mathbb{T}_{h}$ have vertices $\bm{p}_{i}$} ($i=1,\ldots,3$). We assume that $\overline{{{\bm{p}}_{2}{\bm{p}}_{3}}}$ is the longest edge of T{T}, that is, $h_{{T}}:=|{{\bm{p}}_{2}-{\bm{p}}_{3}}|$. We set $h_{1}=|{{\bm{p}}_{1}-{\bm{p}}_{2}}|$ and $h_{2}=|{{\bm{p}}_{1}-{\bm{p}}_{3}}|$. We then assume that $h_{2}\leq h_{1}$. Because $\frac{1}{2}h_{T}<h_{1}\leq h_{T}$, $h_{1}\approx h_{T}$}.
Condition 2.4 (Case in which $d=3$).
Let ${T}\in\mathbb{T}_{h}$ have vertices ${{\bm{p}}}_{i}$ ($i=1,\ldots,4$). Let ${L}_{i}$ ($1\leq i\leq 6$) be the edges of T{T}. We denote by ${L}_{\min}$ the edge of T{T} with the minimum length; that is, $|{L}_{\min}|=\min_{1\leq i\leq 6}|{L}_{i}|$. We set $h_{2}:=|{L}_{\min}|$ and assume that
[…]
Among the four edges sharing an endpoint with ${L}_{\min}$, we consider the longest edge ${L}^{({\min})}_{\max}$. Let ${{\bm{p}}}_{1}$ and ${\bm{p}}_{2}$ be the endpoints of edge ${L}^{({\min})}_{\max}$. Thus, we have
$\displaystyle h_{1}=|{L}^{(\min)}_{\max}|=|{{\bm{p}}_{1}-{\bm{p}}_{2}}|.$
We consider cutting $\mathbb{R}^{3}$ with a plane that contains the midpoint of the edge ${L}^{(\min)}_{\max}$ and is perpendicular to the vector ${\bm{p}}_{1}-{\bm{p}}_{2}$}. Thus, there are two cases.
(Type i)
${\bm{p}}_{3}$ and ${\bm{p}}_{4}$ belong to the same half-space;
(Type ii)
${\bm{p}}_{3}$ and ${\bm{p}}_{4}$ belong to different half-spaces.
In each case, we set
(Type i)
${\bm{p}}_{1}$ and ${\bm{p}}_{3}$ as the endpoints of ${L}_{\min}$, that is, $h_{2}=|{\bm{p}}_{1}-{\bm{p}}_{3}|$;
(Type ii)
${\bm{p}}_{2}$ and ${\bm{p}}_{3}$ as the endpoints of ${L}_{\min}$, that is, $h_{2}=|{\bm{p}}_{2}-{\bm{p}}_{3}|$.
Finally, we set $h_{3}=|{\bm{p}}_{1}-{\bm{p}}_{4}|$. We implicitly assume that ${\bm{p}}_{1}$ and ${\bm{p}}_{4}$ belong to the same half-space. Additionally, note that $h_{1}\approx h_{T}$}.
Lemma 2.5.
It holds that
$\displaystyle\|\widehat{{A}}\|_{2}$ $\leq h_{T},\quad\|\widehat{{A}}\|_{2}\|\widehat{{A}}^{-1}\|_{2}=\frac{\max\{h_{1},\cdots,h_{d}\}}{\min\{h_{1},\cdots,h_{d}\}},$ (2.16a)
$\|\widetilde{{A}}\|_{2}$ $\leq\begin{cases}\sqrt{2}\quad\text{if $d=2$},\\ 2\quad\text{if $d=3$},\end{cases}\quad\|\widetilde{{A}}\|_{2}\|\widetilde{{A}}^{-1}\|_{2}\leq\begin{cases}\frac{h_{1}h_{2}}{|T|_{2}}=\frac{H_{T}}{h_{T}}\quad\text{if $d=2$},\\ \frac{2}{3}\frac{h_{1}h_{2}h_{3}}{|T|_{3}}=\frac{2}{3}\frac{H_{T}}{h_{T}}\quad\text{if $d=3$},\end{cases}$ (2.16b)
$\|{A}_{T}\|_{2}$ $=1,\quad\|{A}_{T}^{-1}\|_{2}=1.$ (2.16c)
where a parameter $H_{T}$ is defined in Definition 3.1. Furthermore, we have
$\displaystyle|\det({A}_{\widetilde{T}})|=|\det(\widetilde{{A}})||\det(\widehat{{A}})|=\frac{|T|_{d}}{|\widetilde{T}|_{d}}\frac{|\widetilde{T}|_{d}}{|\widehat{T}|_{d}}=d!|T|_{d},\quad|\det({A}_{T})|=1,$ (2.17)
where $\|{A}\|_{2}$ denotes the operator norm of A{A}.
Proof.
A proof can be found in [15, Lemma 2]. ∎
2.7 Two-step Piola transforms
We adopt the following two-step Piola transformations.
Definition 2.6 (Two-step Piola transforms).
Let $V(\widehat{T}):=\mathcal{C}(\widehat{T})^{d}$. The Piola transformation $\Psi:={\Psi}_{\widetilde{T}}\circ{\Psi}_{\widehat{T}}:V(\widehat{T})\to V(T)$ is defined as
[…]
with two Piola transformations:
[…]
2.8 Additional notations
We define the vectors ${\bm{r}}_{n}\in\mathbb{R}^{d}$, $n=1,\ldots,d$ as follows: If $d=2$,
$\displaystyle{\bm{r}}_{1}:=\frac{{\bm{p}}_{2}-{\bm{p}}_{1}}{|{\bm{p}}_{2}-{\bm{p}}_{1}|},\quad{\bm{r}}_{2}:=\frac{{\bm{p}}_{3}-{\bm{p}}_{1}}{|{\bm{p}}_{3}-{\bm{p}}_{1}|},$
see Fig. 1, and if $d=3$,
${\bm{r}}_{1}:=\frac{{\bm{p}}_{2}-{\bm{p}}_{1}}{|{\bm{p}}_{2}-{\bm{p}}_{1}|},\quad{\bm{r}}_{3}:=\frac{{\bm{p}}_{4}-{\bm{p}}_{1}}{|{\bm{p}}_{4}-{\bm{p}}_{1}|},\quad\begin{cases}\displaystyle{\bm{r}}_{2}:=\frac{{\bm{p}}_{3}-{\bm{p}}_{1}}{|{\bm{p}}_{3}-{\bm{p}}_{1}|},\quad\text{for (Type i)},\\ \displaystyle{\bm{r}}_{2}:=\frac{{\bm{p}}_{3}-{\bm{p}}_{2}}{|{\bm{p}}_{3}-{\bm{p}}_{2}|}\quad\text{for (Type ii)},\end{cases}$
see Fig 3 for (Type i) and Fig 3 for (Type ii).
Fig. 1: Two-step affine mapping and vectors $r_{i}$, $i=1,2$
Fig. 2: (Type i) Vectors $r_{i}$, $i=1,2,3$
Fig. 3: (Type ii) Vectors $r_{i}$, $i=1,2,3$
For a sufficiently smooth function $\varphi$ and a vector function ${\bm{v}}:=(v_{1},\ldots,v_{d})^{\top}$, we define the directional derivative for $i\in\{1,\ldots,d\}$ as
$\displaystyle\frac{\partial\varphi}{\partial{{\bm{r}}_{i}}}$ $:=({\bm{r}}_{i}\cdot{\bm{\nabla}}_{x})\varphi=\sum_{i_{0}=1}^{d}({\bm{r}}_{i})_{i_{0}}\frac{\partial\varphi}{\partial x_{i_{0}}},$
$\frac{\partial{\bm{v}}}{\partial{\bm{r}}_{i}}$ $:=\left(\frac{\partial v_{1}}{\partial{\bm{r}}_{i}},\ldots,\frac{\partial v_{d}}{\partial{\bm{r}}_{i}}\right)^{\top}=(({\bm{r}}_{i}\cdot{\bm{\nabla}}_{x})v_{1},\ldots,({\bm{r}}_{i}\cdot{\bm{\nabla}}_{x})v_{d})^{\top}.$
For a multiindex $\beta=(\beta_{1},\ldots,\beta_{d})\in\mathbb{N}_{0}^{d}$, we use the notation
$\displaystyle\partial^{\beta}\varphi:=\frac{\partial^{|\beta|}\varphi}{\partial x_{1}^{\beta_{1}}\ldots\partial x_{d}^{\beta_{d}}},\quad\partial^{\beta}_{r}\varphi:=\frac{\partial^{|\beta|}\varphi}{\partial{\bm{r}}_{1}^{\beta_{1}}\ldots\partial{\bm{r}}_{d}^{\beta_{d}}},\quad h^{\beta}:=h_{1}^{\beta_{1}}\cdots h_{d}^{\beta_{d}}.$ (2.19)
We note that $\partial^{\beta}\varphi\neq\partial^{\beta}_{r}\varphi$.
2.9 $L^{2}$-orthogonal projection
This section considers error estimates of the $L^{2}$-orthogonal projection, e.g., for a standard argument, see [4, Section 11.5.3]. Here, the discussion is based on the two-step affine mapping.
[…]
Because $^{\Pi}_{\widehat{T}}^{0}\hat{\varphi}-\hat{\varphi}$ and $^{\Pi}_{\widehat{T}}^{0}\hat{\varphi}-\hat{q}$ are $L^{2}$-orthogonal for any $\hat{q}\in\widehat{P}$, the Pythagorean identity yields
$\displaystyle\|\hat{\varphi}-\hat{q}\|^{2}_{L^{2}(\widehat{T})}=\|\hat{\varphi}-{\Pi}_{\widehat{T}}^{0}\hat{\varphi}\|^{2}_{L^{2}(\widehat{T})}+\|{\Pi}_{\widehat{T}}^{0}\hat{\varphi}-\hat{q}\|^{2}_{L^{2}(\widehat{T})}.$
This implies that
$\displaystyle\Pi_{\widehat{T}}^{0}\hat{\varphi}=\mathop{\mathrm{arg\ min}}_{\hat{q}\in\widehat{P}}\|\hat{\varphi}-\hat{q}\|_{L^{2}(\widehat{T})}$.
[…]
$\displaystyle\widetilde{P}$ $:=\{\hat{q}\circ\Phi_{\widetilde{T}}^{-1};\ \hat{q}\in\widehat{{P}}\}$,
${P}$ $:=\{\tilde{q}\circ\Phi_{T};\ \tilde{q}\in\widetilde{{P}}\}$.
The $L^{2}$-orthognal projections onto $\widehat{P}$ and $P$ are respectively the linear operators $\Pi^{0}_{\widetilde{T}}:L^{1}(\widetilde{T})\to\widetilde{P}$ and $\Pi^{0}_{{T}}:L^{1}({T})\to{P}$ defined as
[…]
$\int_{{T}}({\Pi}_{{T}}^{0}{\varphi}-{\varphi})d{x}=0\quad\forall{\varphi}\in L^{1}({T}).$
Then, $\widetilde{P}$ and $P$ are respectively pointwise invariant under $\Pi_{\widetilde{T}}^{0}$ and $\Pi_{T}^{0}$.
We also define the global interpolation $\Pi_{h}^{0}$ to space $V_{h}^{DC(0)}$ as
$\displaystyle(\Pi_{h}^{0}\varphi)|_{T}:=\Pi_{{T}}^{0}(\varphi|_{T})\quad\forall T\in\mathbb{T}_{h},\quad\forall\varphi\in L^{1}(\Omega).$
Lemma 2.7.
Let $q\in[1,\infty)$. It holds that
[…]
Proof.
Because all the norms in the finite-dimensional space $\widehat{P}$ are equivalent, there exist $\hat{c}_{1}$ and $\hat{c}_{2}$, depending on $\widehat{T}$, such that
$\displaystyle\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q}(\widehat{T})}$ $\leq\hat{c}_{1}\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{2}(\widehat{T})}$, (2.22)
$\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q^{\prime}}(\widehat{T})}$ $\leq\hat{c}_{2}\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q}(\widehat{T})}$, (2.23)
where $\frac{1}{q}+\frac{1}{q^{\prime}}=1$. Then,
$\displaystyle\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q}(\widehat{T})}^{2}$ $\leq c\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{2}(\widehat{T})}^{2}=c\int_{\widehat{T}}\hat{\varphi}\Pi_{\widehat{T}}^{0}\hat{\varphi}d\hat{x}$
$\leq c\|\hat{\varphi}\|_{L^{q}(\widehat{T})}\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q^{\prime}}(\widehat{T})}$
$\leq c\|\hat{\varphi}\|_{L^{q}(\widehat{T})}\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q}(\widehat{T})}$,
where we used (2.22), (2.23), (2.20) with $\hat{q}:=\Pi_{\widehat{T}}^{0}\hat{\varphi}$, and the Hölder’s inequality with $\frac{1}{q}+\frac{1}{q^{\prime}}=1$. This proves the target inequality. ∎
The following theorem gives an anisotropic error estimate of the projection $\Pi_{T}^{0}$.
Theorem 2.8.
Let $p\in[1,\infty)$ and $q\in[1,\infty)$ be such that
$\displaystyle W^{1,p}({T})\hookrightarrow L^{q}({T}),$ (2.24)
[…]
$\displaystyle\|\Pi_{T}^{0}\varphi-\varphi\|_{L^{q}(T)}\leq c|T|_{d}^{\frac{1}{q}-\frac{1}{p}}\sum_{i=1}^{d}h_{i}\left\|\frac{\partial\varphi}{\partial\bm{r}_{i}}\right\|_{L^{p}(T)}.$ (2.25)
Proof.
Using the scaling argument, we have
$\displaystyle\|\Pi_{T}^{0}\varphi-\varphi\|_{L^{q}(T)}$ $=c|\det({A}_{\widetilde{T}})|^{\frac{1}{q}}\|\Pi_{\widehat{T}}^{0}\hat{\varphi}-\hat{\varphi}\|_{L^{q}(\widehat{T})}$. (2.26)
where we used $|\det(A_{T})|=1$. For any $\hat{\eta}\in\mathbb{P}^{0}\subset\widehat{P}$, from the triangle inequality and $\Pi_{\widehat{T}}^{0}\hat{\eta}=\hat{\eta}$, we have
$\displaystyle\|\Pi_{\widehat{T}}^{0}\hat{\varphi}-\hat{\varphi}\|_{L^{q}(\widehat{T})}$ $\leq\|\Pi_{\widehat{T}}^{0}(\hat{\varphi}-\hat{\eta})\|_{L^{q}(\widehat{T})}+\|\hat{\eta}-\hat{\varphi}\|_{L^{q}(\widehat{T})}$. (2.27)
Using (2.21) for the first term on the right-hand side of (2.27), we have
$\displaystyle\|\Pi_{\widehat{T}}^{0}(\hat{\varphi}-\hat{\eta})\|_{L^{q}(\widehat{T})}\leq c\|\hat{\varphi}-\hat{\eta}\|_{L^{q}(\widehat{T})}$. (2.28)
Using the Sobolev embedding theorem for the second term on the right-hand side of (2.27) and (2.28), we obtain
$\displaystyle\|\hat{\varphi}-\hat{\eta}\|_{L^{q}(\widehat{T})}\leq c\|\hat{\varphi}-\hat{\eta}\|_{W^{1,p}(\widehat{T})}$. (2.29)
Combining (2.26), (2.27), (2.28), and (2.29), we have
$\displaystyle\|\Pi_{T}^{0}\varphi-\varphi\|_{L^{q}(T)}$ $\leq c|\det({A}_{\widetilde{T}})|^{\frac{1}{q}}\inf_{\hat{\eta}\in\mathbb{P}^{0}}\|\hat{\varphi}-\hat{\eta}\|_{W^{1,p}(\widehat{T})}$. (2.30)
[…]
$\displaystyle|\hat{\varphi}-\hat{\eta}_{\beta}|_{W^{t,p}(\widehat{T})}\leq C^{BH}(\widehat{T})|\hat{\varphi}|_{W^{1,p}(\widehat{T})},\quad t=0,1.$ (2.31)
Using the new scaling argument [15, Lemma 6] ($\ell=1$ and $m=0$) and (2.31), we then have
$\displaystyle\|\hat{\varphi}-\hat{\eta}_{\beta}\|_{W^{1,p}(\widehat{T})}$ $\leq c|\hat{\varphi}|_{W^{1,p}(\widehat{T})}$
$\leq c|\det({A}_{\widetilde{T}})|^{-\frac{1}{p}}\sum_{i=1}^{d}h_{i}\left\|\frac{\partial\varphi}{\partial\bm{r}_{i}}\right\|_{L^{p}(T)}.$ (2.32)
Therefore, combining (2.30) and (2.32), we have (2.25). Here, we used
$\displaystyle|\det({A}_{\widetilde{T}})|=|\det(\widetilde{{A}})||\det(\widehat{{A}})|=\frac{|T|_{d}}{|\widetilde{T}|_{d}}\frac{|\widetilde{T}|_{d}}{|\widehat{T}|_{d}}=d!|T|_{d}.$
3 Semi-regular geometric mesh condition
3.1 New geometric mesh condition
We proposed a new geometric parameter $H_{T}$ in [13, 15]. This parameter represents the flatness of a simplex.
Definition 3.1.
We define the parameter $H_{{T}}$ as
$\displaystyle H_{{T}}:=\frac{\prod_{i=1}^{d}h_{i}}{|{T}|_{d}}h_{{T}}.$
The following geometric condition is equivalent to the maximum angle condition ([12]).
Assumption 1.
A family of meshes $\{\mathbb{T}_{h}\}$ has a semi-regular property if there exists $\gamma_{0}\mathop{\textgreater}0$ such that
$\displaystyle\frac{H_{T}}{h_{T}}\leq\gamma_{0}\quad\forall\mathbb{T}_{h}\in\{\mathbb{T}_{h}\},\quad\forall T\in\mathbb{T}_{h}.$ (3.1)
The quantity $H_{T}/h_{T}$ can be easily calculated in the numerical process of finite element methods. Therefore, the new condition may be useful in the case of adaptive finite element methods. We expect the new mesh condition to become an alternative to the maximum-angle condition.
3.2 RT finite element interpolation operator
For $T\in\mathbb{T}_{h}$, the local RT polynomial space is defined as
$\displaystyle\mathbb{RT}^{0}(T):=\mathbb{P}^{0}(T)^{d}+{\bm{x}}\mathbb{P}^{0}(T),\quad{\bm{x}}\in\mathbb{R}^{d}.$ (3.2)
Let ${I}_{T}^{RT}:W^{1,1}(T)^{d}\to\mathbb{RT}^{0}(T)$ be the RT interpolation operator such that for any ${\bm{v}}\in W^{1,1}(T)^{d}$,
$\displaystyle{I}_{T}^{RT}:W^{1,1}(T)^{d}\ni{\bm{v}}\mapsto{I}_{T}^{RT}{\bm{v}}:=\sum_{i=1}^{d+1}\left(\int_{{F}_{T,i}}{\bm{v}}\cdot{\bm{n}}_{T,i}d{s}\right){\bm{\theta}}_{T,i}^{RT}\in\mathbb{RT}^{0}(T),$ (3.3)
where ${\bm{\theta}}_{T,i}^{RT}$ is the local shape basis function (e.g. [4, p. 162]) and ${\bm{n}}_{T,i}$ is a fixed unit normal to ${F}_{T,i}$.
The RT finite-element space is defined as follows:
$\displaystyle V^{RT}_{h}$ $:=\{{\bm{v}_{h}}\in L^{1}(\Omega)^{d}:\ {\bm{v}_{h}}|_{T}\in\mathbb{RT}^{0}(T),\ \forall T\in\mathbb{T}_{h},\ [\![{\bm{v}_{h}}\cdot{\bm{n}}]\!]_{F}=0,\ \forall F\in\mathcal{F}_{h}^{i}\}.$
We define the following global RT interpolation ${I}_{h}^{RT}:W^{1,1}(\Omega)^{d}\to V^{RT}_{h}$ as
$\displaystyle({I}_{h}^{RT}{\bm{v}})|_{T}:={I}_{T}^{RT}({\bm{v}}|_{T})\quad\forall T\in\mathbb{T}_{h},\quad\forall{\bm{v}}\in W^{1,1}(\Omega)^{d}.$
The following two lemmata are divided into the element on $\mathfrak{T}^{(2)}$ or $\mathfrak{T}_{1}^{(3)}$ and the element on $\mathfrak{T}_{2}^{(3)}$ in Section 2.6.
Lemma 3.2.
[…]
$\displaystyle\|{I_{T}^{RT}\bm{v}}\|_{L^{p}({T})^{d}}$ $\leq c\left[\frac{H_{T}}{h_{T}}\left(\|\bm{v}\|_{L^{p}(T)^{d}}+\sum_{|\varepsilon|=1}h^{\varepsilon}\left\|\partial_{\bm{r}}^{\varepsilon}\bm{v}\right\|_{L^{p}(T)^{d}}\right)+h_{T}\|\bm{\nabla}\cdot\bm{v}\|_{L^{p}({T})}\right].$ (3.4)
Proof.
A proof can be found in [6, Lemma 8]. ∎
Lemma 3.3.
[…]
$\displaystyle\|{I_{T}^{RT}\bm{v}}\|_{L^{p}({T})^{3}}$ $\leq c\frac{H_{T}}{h_{T}}\left[\|\bm{v}\|_{L^{p}(T)^{3}}+h_{T}\sum_{k=1}^{3}\left\|\frac{\partial\bm{v}}{\partial\bm{r}_{k}}\right\|_{L^{p}(T)^{3}}\right].$ (3.5)
Proof.
A proof can be found in [6, Lemma 8]. ∎
Lemmata 3.2 and 3.3 yield the following corollary.
Corollary 3.4 (Stability).
Let $p\in[1,\infty)$. We impose Assumption 1 with $h\leq 1$. Then,
$\displaystyle\|{I_{h}^{RT}\bm{v}}\|_{L^{p}({\Omega})^{d}}$ $\leq C(\gamma_{0})\|\bm{v}\|_{W^{1,p}(\Omega)^{d}}\quad\forall\bm{v}\in W^{1,p}(\Omega)^{d}.$
The following two theorems are based on an element $T$ satisfying Type i or Type ii in Section 2.6 when $d=3$.
Theorem 3.5.
[…]
$\displaystyle\|{I}_{T}^{RT}\bm{v}-\bm{v}\|_{L^{p}(T)^{d}}$ $\leq c\left(\frac{H_{T}}{h_{T}}\sum_{i=1}^{d}h_{i}\left\|\frac{\partial\bm{v}}{\partial r_{i}}\right\|_{L^{p}(T)^{d}}+h_{T}\|\mathop{\mathrm{div}}\bm{v}\|_{L^{p}({T})}\right).$ (3.6)
Proof.
A proof can be found in [6, Theorem 2]. ∎
Theorem 3.6.
[…]
$\|{I}_{T}^{RT}\bm{v}-\bm{v}\|_{L^{p}(T)^{3}}\leq c\frac{H_{T}}{h_{T}}\Biggl{(}h_{T}|\bm{v}|_{W^{1,p}(T)^{3}}\Biggr{)}.$ (3.7)
Proof.
A proof can be found in [6, Theorem 3]. ∎
Corollary 3.7.
Let $p\in[1,\infty)$. We impose Assumption 1. Then, for any $\hat{\bm{v}}\in W^{1,p}(\widehat{T})^{d}$ with $\tilde{\bm{v}}={\Psi}_{\widehat{T}}\hat{\bm{v}}$ and ${\bm{v}}={\Psi}_{\widetilde{T}}\tilde{\bm{v}}$,
$\displaystyle\|{I}_{T}^{RT}\bm{v}-\bm{v}\|_{L^{p}(T)^{d}}$ $\leq C_{RT}(\gamma_{0})h_{T}|\bm{v}|_{W^{1,p}(T)^{d}}.$ (3.8)
4 Lemmata for analysis
Lemma 4.1 (Bogovskiĭ-type lemma).
Let $D\subset\mathbb{R}^{d}$, $d\in\{2,3\}$ be a bounded, connected domain that is a finite union of open sets with Lipschitz-continuous boundaries. Then, there exists an operator $\mathcal{L}$ from $L^{1}_{0}(D)$ into $L^{1}(D)^{d}$ such that
$\displaystyle\mathop{\mathrm{div}}(\mathcal{L}f)=f\quad\forall f\in L^{1}_{0}(D).$ (4.1)
Furthermore, the operator $\mathcal{L}$ maps $\mathcal{D}(D)\cap L^{1}_{0}(D)$ into $\mathcal{D}(D)^{d}$ and for each real $r\in(1,\infty)$, is continuous from $L^{r}_{0}(D)$ into $W^{1,r}_{0}(D)^{d}$; there exists a constant $C_{r}$ such that
$\displaystyle|\mathcal{L}f|_{W^{1,r}(D)^{d}}\leq C_{r}\|f\|_{L^{r}(D)}\quad\forall f\in L^{r}_{0}(D),$ (4.2)
where $L^{r}_{0}(D):=\{v\in L^{r}(D):\ \int_{D}vdx=0\}$.
Proof.
A proof is available in [5]; see also [1, Theorem 1.4.5] for further reference. ∎
The subsequent relation is integral to our analysis of anisotropic meshes.
Lemma 4.2.
For any $\bm{\tau}_{h}\in V_{h}^{RT}$ and $\psi_{h}\in V_{h}^{DC(1)}$,
$\int_{\Omega}\left(\bm{\tau}_{h}\cdot\bm{\nabla}_{h}\psi_{h}+\mathop{\mathrm{div}}\bm{\tau}_{h}\psi_{h}\right)dx$
$\quad=\sum_{F\in\mathcal{F}_{h}^{i}}\int_{F}\{\!\{\bm{\tau}_{h}\}\!\}_{\omega,F}\cdot\bm{n}_{F}\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}ds+\sum_{F\in\mathcal{F}_{h}^{\partial}}\int_{F}(\bm{\tau}_{h}\cdot\bm{n}_{F})\Pi_{F}^{0}\psi_{h}ds.$ (4.3)
Proof.
A proof can be found in [7, Lemma 3]. ∎
Corollary 4.3.
Let $p^{\prime}\in[1,\infty)$. For any $\bm{w}\in W^{1,p^{\prime}}(\Omega)^{d}$ and $\psi_{h}\in V_{h}^{DC(1)}$,
$\int_{\Omega}\left({I}_{h}^{RT}\bm{w}\cdot\bm{\nabla}_{h}\psi_{h}+\mathop{\mathrm{div}}{I}_{h}^{RT}\bm{w}\psi_{h}\right)dx$
$\quad=\sum_{F\in\mathcal{F}_{h}^{i}}\int_{F}\{\!\{\bm{w}\}\!\}_{\omega,F}\cdot\bm{n}_{F}\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}ds+\sum_{F\in\mathcal{F}_{h}^{\partial}}\int_{F}(\bm{w}\cdot\bm{n}_{F})\Pi_{F}^{0}\psi_{h}ds.$ (4.4)
Proof.
Let $p^{\prime}\in[1,\infty)$. For any $\bm{w}\in W^{1,p^{\prime}}(\Omega)^{d}$, we set $\bm{\tau}_{h}:={I}_{h}^{RT}\bm{w}$ in (4.3). Using the definition of $I_{h}^{RT}$ yields (4.4). ∎
Lemma 4.4.
Let $p^{\prime}\in(1,\infty)$ and $h\leq 1$. For any $\bm{w}\in W^{1,p^{\prime}}(\Omega)^{d}$ and $\psi_{h}\in V_{h}^{DC(1)}$,
$\displaystyle\left|\sum_{F\in\mathcal{F}_{h}^{i}}\int_{F}\{\!\{\bm{w}\}\!\}_{\omega,F}\cdot\bm{n}_{F}\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}ds\right|$ $\leq c|\psi_{h}|_{p,J}\|\bm{w}\|_{W^{1,p^{\prime}}(\Omega)^{d}},$ (4.5)
$\left|\sum_{F\in\mathcal{F}_{h}^{\partial}}\int_{F}(\bm{w}\cdot\bm{n}_{F})\Pi_{F}^{0}\psi_{h}ds\right|$ $\leq c|\psi_{h}|_{p,J}\|\bm{w}\|_{W^{1,p^{\prime}}(\Omega)^{d}},$ (4.6)
where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$.
Proof.
Suppose that $F\in\mathcal{F}_{h}^{i}$ with $F=T_{+}\cap T_{-}$, $T_{+},T_{-}\in\mathbb{T}_{h}$. Using the Hölder’s inequality, the weighted average and the trace inequality (2.1) yields
$\int_{F}\left|\{\!\{\bm{w}\}\!\}_{\omega,F}\cdot\bm{n}_{F}\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\right|ds$
$\quad\leq\|\omega_{T_{+},F}{\bm{w}_{+}}+\omega_{T_{-},F}{\bm{w}_{-}}\|_{L^{p^{\prime}}(F)^{d}}\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}$
$\quad\leq\left(\omega_{T_{+},F}\|\bm{w}_{+}\|_{L^{p^{\prime}}(F)^{d}}+\omega_{T_{-},F}\|\bm{w}_{-}\|_{L^{p^{\prime}}(F)^{d}}\right)\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}$
$\quad\leq c\left(\|\bm{w}_{+}\|_{L^{p^{\prime}}(T_{+})^{d}}+h_{T_{+}}^{\frac{1}{p^{\prime}}}\|\bm{w}_{+}\|_{L^{p^{\prime}}(T_{+})^{d}}^{1-\frac{1}{p^{\prime}}}|\bm{w}_{+}|_{W^{1,p^{\prime}}(T_{+})^{d}}^{\frac{1}{p^{\prime}}}\right)\omega_{T_{+},F}\ell_{T_{+},F}^{-\frac{1}{p^{\prime}}}\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}$
$\quad\quad+c\left(\|\bm{w}_{-}\|_{L^{p^{\prime}}(T_{-})^{d}}+h_{T_{-}}^{\frac{1}{p^{\prime}}}\|\bm{w}_{-}\|_{L^{p^{\prime}}(T_{-})^{d}}^{1-\frac{1}{p^{\prime}}}|\bm{w}_{-}|_{W^{1,p^{\prime}}(T_{-})^{d}}^{\frac{1}{p^{\prime}}}\right)\omega_{T_{-},F}\ell_{T_{-},F}^{-\frac{1}{p^{\prime}}}\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}$
$\quad\leq c\Bigg{\{}\left(\|\bm{w}_{+}\|_{L^{p^{\prime}}(T_{+})^{d}}+h_{T_{+}}^{\frac{1}{p^{\prime}}}\|\bm{w}_{+}\|_{L^{p^{\prime}}(T_{+})^{d}}^{1-\frac{1}{p^{\prime}}}|\bm{w}_{+}|_{W^{1,p^{\prime}}(T_{+})^{d}}^{\frac{1}{p^{\prime}}}\right)^{p^{\prime}}$
$\quad\quad+\left(\|\bm{w}_{-}\|_{L^{p^{\prime}}(T_{-})^{d}}+h_{T_{-}}^{\frac{1}{p^{\prime}}}\|\bm{w}_{-}\|_{L^{p^{\prime}}(T_{-})^{d}}^{1-\frac{1}{p^{\prime}}}|\bm{w}_{-}|_{W^{1,p^{\prime}}(T_{-})^{d}}^{\frac{1}{p^{\prime}}}\right)^{p^{\prime}}\Bigg{\}}^{\frac{1}{p^{\prime}}}$
$\quad\quad\times\left(\omega_{T_{+},F}^{p}\ell_{T_{+},F}^{-\frac{p}{p^{\prime}}}+\omega_{T_{-},F}^{p}\ell_{T_{-},F}^{-\frac{p}{p^{\prime}}}\right)^{\frac{1}{p}}\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}.$
Using the Hölder’s inequality again, we have
$\left|\sum_{F\in\mathcal{F}_{h}^{i}}\int_{F}\{\!\{\bm{w}\}\!\}_{\omega,F}\cdot\bm{n}_{F}\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}ds\right|$
$\quad\leq c\sum_{F\in\mathcal{F}_{h}^{i}}\left(\omega_{T_{+},F}^{p}\ell_{T_{+},F}^{-\frac{p}{p^{\prime}}}+\omega_{T_{-},F}^{p}\ell_{T_{-},F}^{-\frac{p}{p^{\prime}}}\right)^{\frac{1}{p}}\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}$
$\quad\quad\times\sum_{T\in\mathbb{T}_{F}}\left(\|\bm{w}\|_{L^{p^{\prime}}(T)^{d}}+h_{T}^{\frac{1}{p^{\prime}}}\|\bm{w}\|_{L^{p^{\prime}}(T)^{d}}^{1-\frac{1}{p^{\prime}}}|\bm{w}|_{W^{1,p^{\prime}}(T)^{d}}^{\frac{1}{p^{\prime}}}\right)$
$\quad\leq c\left(\sum_{F\in\mathcal{F}_{h}^{i}}\left(\omega_{T_{+},F}^{p}\ell_{T_{+},F}^{-\frac{p}{p^{\prime}}}+\omega_{T_{-},F}^{p}\ell_{T_{-},F}^{-\frac{p}{p^{\prime}}}\right)\|\Pi_{F}^{0}[\![\psi_{h}]\!]_{F}\|_{L^{p}(F)}^{p}\right)^{\frac{1}{p}}$
$\quad\quad\times\left(\sum_{F\in\mathcal{F}_{h}^{i}}\sum_{T\in\mathbb{T}_{F}}\left(\|\bm{w}\|_{L^{p^{\prime}}(T)^{d}}+h_{T}^{\frac{1}{p^{\prime}}}\|\bm{w}\|_{L^{p^{\prime}}(T)^{d}}^{1-\frac{1}{p^{\prime}}}|\bm{w}|_{W^{1,p^{\prime}}(T)^{d}}^{\frac{1}{p^{\prime}}}\right)^{p^{\prime}}\right)^{\frac{1}{p^{\prime}}}$
$\quad\leq c|\psi_{h}|_{p,J}\|\bm{w}\|_{W^{1,p^{\prime}}(\Omega)^{d}},$
which is the inequality (4.5) together with the weight (2.9) and Jensen inequality. Here, $\mathbb{T}_{F}$ denotes the set of the simplices in $\mathbb{T}_{h}$ that share $F$ as a common face.
By an analogous argument, the estimate (4.6) holds. ∎
5 Main theorem
In this section, we present the discrete Sobolev inequalities on anisotropic meshes. Our goal is to prove the discrete Sobolev inequalities under the semi-regular mesh condition. The proof combines an anisotropic trace estimate, two-step affine/Piola maps, RT-interpolation stability, and a weighted discrete integration-by-parts formula.
Let $q\in(1,\infty)$ with $\frac{1}{q}+\frac{1}{q^{\prime}}=1$. For any $u\in L^{q}(\Omega)$ and $v\in L^{q^{\prime}}(\Omega)$,
$\displaystyle\langle u,v\rangle:=\int_{\Omega}uvdx=\sum_{T\in\mathbb{T}_{h}}\int_{T}uvdx.$
We define a functional space as
$\displaystyle X_{0}$ $:=\left\{\psi\in L^{q^{\prime}}(\Omega):\ \|\psi\|_{L^{q^{\prime}}(\Omega)}=1\right\}.$
Theorem 5.1 (Discrete $L^{q}$-$L^{p}$ Sobolev inequality).
Let $h\leq 1$. Suppose that Assumption 1 holds. Let $p\in(1,\infty)$ and $q\in(1,\infty)$ be such that
$\displaystyle W^{1,p}({T})\hookrightarrow L^{q}({T}),$ (5.1)
that is $1-\frac{d}{p}\geq-\frac{d}{q}$ and $q\leq p$. Then,
$\displaystyle\|\varphi_{h}\|_{L^{q}(\Omega)}\leq C^{dS0(q,p)}|\varphi_{h}|_{p,V_{h}}\quad\forall\varphi_{h}\in V_{h}\in\{V_{h}^{DCCR},V_{h}^{CR},V_{h0}^{CR}\},$ (5.2)
where $C^{dS0(q,p)}$ is a positive constant independent of $h$ and $\varphi_{h}$.
Proof.
Let $\varphi_{h}\in V_{h}$. Using the triangle inequality yields
$\displaystyle\|\varphi_{h}\|_{L^{q}(\Omega)}\leq\|\varphi_{h}-\Pi_{h}^{0}\varphi_{h}\|_{L^{q}(\Omega)}+\|\Pi_{h}^{0}\varphi_{h}\|_{L^{q}(\Omega)}.$ (5.3)
By using the estimate (2.25) for each element, we obtain the following
$\displaystyle\|\varphi_{h}-\Pi_{h}^{0}\varphi_{h}\|_{L^{q}(\Omega)}$ $\leq c|\varphi_{h}|_{W^{1,p}(\mathbb{T}_{h})}.$ (5.4)
Because $\Omega$ is bounded and $q\leq p$, then $L^{p}(\Omega)\hookrightarrow L^{q}(\Omega)$ and
$\displaystyle\|\Pi_{h}^{0}\varphi_{h}\|_{L^{q}(\Omega)}$ $\leq C(\Omega)\|\Pi_{h}^{0}\varphi_{h}\|_{L^{p}(\Omega)}.$ (5.5)
The $L^{p}$-norm of $\Pi_{h}^{0}\varphi_{h}$ can be written in dual form
$\displaystyle\|\Pi_{h}^{0}\varphi_{h}\|_{L^{p}(\Omega)}=\sup_{\psi\in X_{0}}\langle\Pi_{h}^{0}\varphi_{h},\psi\rangle.$
For any $\psi\in X_{0}$, we set $\psi_{h}:=\Pi_{h}^{0}\psi$. Then,
$\displaystyle\psi_{h}|_{T}=\frac{1}{|T|_{d}}\int_{T}\psi dx.$
Therefore, we have
$\displaystyle\langle\Pi_{h}^{0}\varphi_{h},\psi\rangle$ $=\sum_{T\in\mathbb{T}_{h}}\Pi_{T}^{0}\varphi_{h}\int_{T}\psi dx=\sum_{T\in\mathbb{T}_{h}}\Pi_{T}^{0}\varphi_{h}\psi_{h}|_{T}|T|_{d}.$
On the other hand,
$\displaystyle\langle\Pi_{h}^{0}\varphi_{h},\psi_{h}\rangle=\sum_{T\in\mathbb{T}_{h}}\int_{T}\Pi_{h}^{0}\varphi_{h}\psi_{h}dx=\sum_{T\in\mathbb{T}_{h}}\Pi_{T}^{0}\varphi_{h}\int_{T}\psi_{h}|_{T}dx=\sum_{T\in\mathbb{T}_{h}}\Pi_{T}^{0}\varphi_{h}\psi_{h}|_{T}|T|_{d},$
which leads to $\langle\Pi_{h}^{0}\varphi_{h},\psi\rangle=\langle\Pi_{h}^{0}\varphi_{h},\psi_{h}\rangle$. Using the Hölder’s inequality yields
$\displaystyle\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}^{p^{\prime}}$ $=\sum_{T\in\mathbb{T}_{h}}\int_{T}|\psi_{h}|^{p^{\prime}}dx=\sum_{T\in\mathbb{T}_{h}}|T|_{d}^{1-p^{\prime}}\left|\int_{T}\psi dx\right|^{p^{\prime}}$
$\leq\sum_{T\in\mathbb{T}_{h}}|T|_{d}^{1-p^{\prime}+\frac{p^{\prime}}{p}}\int_{T}|\psi|^{p^{\prime}}dx=\sum_{T\in\mathbb{T}_{h}}\int_{T}|\psi|^{p^{\prime}}dx=\|\psi\|_{L^{p^{\prime}}(\Omega)}^{p^{\prime}},$
which leads to $\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}\leq 1$.
We set $\xi_{h}:=\psi_{h}-\bar{\psi}_{h}$, where $\bar{\psi}_{h}:=\frac{1}{|\Omega|_{d}}\int_{\Omega}\psi_{h}dx$. Then, $\int_{\Omega}\xi_{h}dx=0$. From the Bogovskiĭ-type lemma (Lemma 4.1), there exists $\bm{v}_{0}\in W_{0}^{1,p^{\prime}}(\Omega)^{d}$ such that
$\displaystyle\mathop{\mathrm{div}}\bm{v}_{0}$ $=\xi_{h},$
$|\bm{v}_{0}|_{W^{1,p^{\prime}}(\Omega)^{d}}$ $\leq C_{p^{\prime}}\|\xi_{h}\|_{L^{p^{\prime}}(\Omega)}.$
Setting $\bm{v}_{g}(\bm{x}):=\frac{\bar{\psi}_{h}}{d}(\bm{x}-\bm{x}_{g})$, where $\bm{x}_{g}:=\frac{1}{|\Omega|}\int_{\Omega}\bm{x}dx$, $\bm{x}\in\Omega$. Then, $\bm{\nabla}\bm{v}_{g}=\frac{\bar{\psi}_{h}}{d}I_{d}$, where $I_{d}$ is the $d\times d$ unit matrix. This implies
$\displaystyle\mathop{\mathrm{div}}\bm{v}_{g}=\bar{\psi}_{h}.$
Furthermore, we have
$\displaystyle\|\bm{v}_{g}\|_{W^{1,p^{\prime}}(\Omega)^{d}}^{p^{\prime}}$ $=\|\bm{v}_{g}\|_{L^{p^{\prime}}(\Omega)^{d}}^{p^{\prime}}+\|\bm{\nabla}\bm{v}_{g}\|_{L^{p^{\prime}}(\Omega)^{d\times d}}^{p^{\prime}}$
$\leq\left(\frac{|\bar{\psi}_{h}|}{d}\right)^{p^{\prime}}\mathop{\mathrm{diam}}(\Omega)^{p^{\prime}}|\Omega|_{d}+\left(\frac{|\bar{\psi}_{h}|}{d}\right)^{p^{\prime}}\|I_{d}\|_{F}^{p^{\prime}}|\Omega|_{d}$
$\leq d^{-p^{\prime}}\left(\mathop{\mathrm{diam}}(\Omega)^{p^{\prime}}+d^{\frac{p^{\prime}}{2}}\right)\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}^{p^{\prime}},$
where $\|\cdot\|_{F}$ is the Frobenius norm and we used that $\|I_{d}\|_{F}=d^{\frac{1}{2}}$ and
$\displaystyle|\bar{\psi}_{h}|$ $\leq\frac{1}{|\Omega|_{d}}\int_{\Omega}|\psi_{h}|dx\leq|\Omega|_{d}^{-\frac{1}{p^{\prime}}}\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}.$
We set $\bm{v}:=\bm{v}_{0}+\bm{v}_{g}\in W^{1,p^{\prime}}(\Omega)^{d}$. Then, it holds that
$\displaystyle\mathop{\mathrm{div}}\bm{v}$ $=\mathop{\mathrm{div}}\bm{v}_{0}+\mathop{\mathrm{div}}\bm{v}_{g}=\psi_{h},$
$\|\bm{v}\|_{W^{1,p^{\prime}}(\Omega)^{d}}$ $\leq\|\bm{v}_{0}\|_{W^{1,p^{\prime}}(\Omega)^{d}}+\|\bm{v}_{g}\|_{W^{1,p^{\prime}}(\Omega)^{d}}$
$\leq C(p^{\prime},d,\Omega)\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}\leq c,$
where we used that
$\displaystyle\|\bm{v}_{0}\|_{W^{1,p^{\prime}}(\Omega)^{d}}$ $\leq C(p^{\prime},\Omega)\|\xi_{h}\|_{L^{p^{\prime}}(\Omega)}\leq C(p^{\prime},\Omega)\left(\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}+\|\bar{\psi}_{h}\|_{L^{p^{\prime}}(\Omega)}\right)$
$\leq C(p^{\prime},d,\Omega)\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}.$
It holds that
$\displaystyle\mathop{\mathrm{div}}I_{h}^{RT}\bm{v}=\Pi_{h}^{0}\mathop{\mathrm{div}}\bm{v}=\Pi_{h}^{0}\psi_{h}=\psi_{h}.$
Using the stability of the RT interpolation (Corollary 3.4) yields
$\displaystyle\|I_{h}^{RT}\bm{v}\|_{L^{p^{\prime}}(\Omega)^{d}}$ $\leq c\|\bm{v}\|_{W^{1,p^{\prime}}(\Omega)^{d}}\leq c\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}\leq c.$
From Lemma 4.3 and the definition of $\Pi_{h}^{0}$, we obtain
$\displaystyle\langle\Pi_{h}^{0}\varphi_{h},\psi\rangle$ $=\langle\Pi_{h}^{0}\varphi_{h},\psi_{h}\rangle=\sum_{T\in\mathbb{T}_{h}}\int_{T}\varphi_{h}\psi_{h}dx$
$=\sum_{T\in\mathbb{T}_{h}}\int_{T}\varphi_{h}\mathop{\mathrm{div}}I_{h}^{RT}\bm{v}dx$
$=-\sum_{T\in\mathbb{T}_{h}}\int_{T}{I}_{h}^{RT}\bm{v}\cdot\bm{\nabla}\varphi_{h}dx+\sum_{F\in\mathcal{F}_{h}^{i}}\int_{F}\{\!\{\bm{v}\}\!\}_{\omega,F}\cdot\bm{n}_{F}\Pi_{F}^{0}[\![\varphi_{h}]\!]_{F}ds+\sum_{F\in\mathcal{F}_{h}^{\partial}}\int_{F}(\bm{v}\cdot\bm{n}_{F})\Pi_{F}^{0}\varphi_{h}ds.$
Using Lemma 4.4 and the above results yields
$\displaystyle\left|\langle\Pi_{h}^{0}\varphi_{h},\psi\rangle\right|$ $\leq c|\varphi_{h}|_{W^{1,p}(\mathbb{T}_{h})}+c|\varphi_{h}|_{p,J}\leq c|\varphi_{h}|_{p,V_{h}},$
which leads to
$\displaystyle\|\Pi_{h}^{0}\varphi_{h}\|_{L^{p}(\Omega)}$ $\leq c|\varphi_{h}|_{p,V_{h}}.$ (5.6)
Gathering the inequalities (5.3), (5.4), (5.5) and (5.6), we conclude the target estimate. ∎
Remark 5.2.
When $q\mathop{\textgreater}p$, the proof of Theorem 5.1 can not be applied. The $L^{q}$-norm of $\Pi_{h}^{0}\varphi_{h}$ can be written in dual form
$\displaystyle\|\Pi_{h}^{0}\varphi_{h}\|_{L^{q}(\Omega)}=\sup_{\psi\in X_{0}}\langle\varphi_{h},\psi_{h}\rangle.$
From the bilateral estimate,
$\displaystyle\|\Pi_{h}^{0}\varphi_{h}\|_{L^{q}(\Omega)}$ $\leq\|\varphi_{h}\|_{L^{p}(\Omega)}\sup_{\psi\in X_{0}}\|\psi_{h}\|_{L^{p^{\prime}}(\Omega)}.$
Because $\psi_{h}|_{T}$ is a constant,
$\displaystyle\|\psi_{h}\|_{L^{p^{\prime}}(T)}$ $\leq c|T|_{d}^{\frac{1}{p^{\prime}}-\frac{1}{q^{\prime}}}\|\psi_{h}\|_{L^{q^{\prime}}(T)}\leq c|T|_{d}^{\frac{1}{q}-\frac{1}{p}}\|\psi_{h}\|_{L^{q^{\prime}}(T)}.$
When an isotropic mesh is used, it is expressed as $|T|_{d}\approx h_{T}^{d}$. Then,
$\displaystyle\|\psi_{h}\|_{L^{p^{\prime}}(T)}$ $\leq ch_{T}^{d\left(\frac{1}{q}-\frac{1}{p}\right)}\|\psi_{h}\|_{L^{q^{\prime}}(T)}\to\infty\quad\text{as $h_{T}\to 0$}.$
6 Concluding remarks
This paper presents discrete $L^{q}-L^{p}$ Sobolev inequalities for nonconforming finite elements under a semi-regular mesh condition, with constants that remain independent of the angle and aspect ratio of simplices. The proof integrates anisotropy-aware trace/projection estimates with RT interpolation and face-weighted flux control. These findings extend the classical shape-regular theory to anisotropic partitions and directly support stability and a priori analyses for CR and Nitsche schemes.
However, when $q\mathop{\textgreater}p$, it remains to show the following restricted discrete Sobolev inequality, see [10, Lemma 4]. For that purpose, we imposed a weak elliptic regularity assumption to obtain a discrete Sobolev inequality: Let $p\in[2,\infty)$ if $d=2$, or $p\in[2,6]$ if $d=3$, then $p^{\prime}\in(\frac{2d}{d+2},2]$. We assume that, for any $g\in L^{p^{\prime}}(\Omega)$, the variational problem
$\displaystyle\int_{\Omega}\bm{\nabla}z\cdot\bm{\nabla}wdx=\int_{\Omega}gwdx\quad\forall w\in H_{0}^{1}(\Omega)$
has a unique solution $z\in H_{0}^{1}(\Omega)$ that belongs to $W^{2,p^{\prime}}(\Omega)$ and the elliptic regularity estimate
$\displaystyle\|z\|_{W^{2,p^{\prime}}(\Omega)}\leq c\|g\|_{L^{p^{\prime}}(\Omega)}$
holds. In addition to the weak elliptic regularity assumption, we impose that Assumptions 1 and there exists a positive constant $C^{dS0}$ independent of $T$ and $h$ such that
$\displaystyle\max_{T\in\mathbb{T}_{h}}\left({|{T}|_{d}}^{\frac{1}{p}-\frac{1}{2}}h_{T}\right)\leq C^{(p,2)}.$
Then,
$\displaystyle\|\varphi_{h}\|_{L^{p}(\Omega)}\leq c|\varphi_{h}|_{H^{1}(\mathbb{T}_{h})}\quad\forall\varphi_{h}\in V_{h0}^{CR}.$
In this statement, when $p=2$, the domain is required to be convex to satisfy the weak elliptic regularity assumption. However, in Theorem 5.1, the assumption that the domain is convex is removed. This means that the weak elliptic regularity assumption may be able to be removed. This issue is left for future work.